Which of the following could not be a value of x when the two side lengths of the triangle are 18 and 9?
A. 10
B. 9
C. 10.5
D. 9.5

Answer by JBarnum(2146)

um where is x on the diagram? is there a picture?
from the answer choices im guessing x isnt an agle and isnt the largest side
a^2+b^2=c^2
81+b^2=324
b^2=243
b=15.5885
have noidea what your trying to figure out please make sure the whole problem is listed in question